OpenStudy (anonymous):

Check! If I'm wrong, do not tell me the answer. Trey is biking at a speed of 300 yd/min. What is Trey’s speed in ft/h? 15 ft/h 480 ft/h 6000 ft/h <----- 54,000 ft/h

OpenStudy (anonymous):

None.

OpenStudy (triciaal):

you need to know the conversion factors how many feet make a yard and how many minutes in an hour

OpenStudy (triciaal):

no try again

OpenStudy (anonymous):

\(1 \text{ yd} = 3\text{ ft}\) and \(60 ~\min = 1\text{ hr}\) So you have: \[\dfrac{300~yd}{min}\cdot\dfrac{60~min}{hr}\cdot\dfrac{3~ft}{yd} =\dfrac{300~\cancel{yd}}{\cancel{min}}\cdot\dfrac{60~\cancel{min}}{hr}\cdot\dfrac{3~ft}{\cancel{yd}} =\cdots~?\]

OpenStudy (triciaal):

@CrazyCountryGirl it is usually easier to do the conversion in 2 steps 1) 300 yards / min = how many feet/ min? changing yds to feet 2) this many feet /min = same many feet in how many hours? changing minutes to hours

OpenStudy (triciaal):

the conversion factors given by @micahwood50 are correct

OpenStudy (anonymous):

You cool? @CrazyCountryGirl

OpenStudy (triciaal):

show/post what you have we will show your error then you can do another one to practice

OpenStudy (anonymous):

I'm sorry. I couldn't figure it out with what you guys were saying. I got my sisters help and now I understand it. Thanks!!